How many different license plates with repetition involving 3 letters and 4 digits if 3 letters must appear together at the beginning or at the end of plate.
Found 2 solutions by greenestamps, rothauserc:
Answer by greenestamps(13200)



Number of sequences (permutations) of 3 letters with repetition (assuming the English language alphabet with 26 letters): 26*26*26 = 26^3

Number of permutations of 4 digits with repetition: 10^4

Number of permutations of the group of 3 letters and the group of 4 digits: 2

Number of license plates: (2)(26^3)(10^4)

Use a calculator if the answer is required in a simplified form.


Answer by rothauserc(4718)

There are 26 letters(A-Z) and 10 digits(0-9)
:
Since letters can repeat, we have 26^3 permutations to form letters on the license plate
:
Since the digits can repeat, we have 10^4 permutations to form digits on the license plate
:
Since the letters can appear at the front or the back, we multiply by 2
:
Therefore,
:
There are 2 * 26^3 * 10^4 = 351,520,000 license plates
: